OpenStudy (anonymous):

show that the point is on the unit circle (12/13,-5/13)

OpenStudy (anonymous):

\[\left( \frac{ 12 }{ 13 },-\frac{ 5 }{ 13 } \right)\]

OpenStudy (anonymous):

 and if you could refresh how this works